2) Calculate the amount of heat necessary to raise the temperature of 84.0 g of water from
12.4C to 62.0C. The specific heat of water = 4.18 J/g·C.
1
Expert's answer
2020-03-23T10:14:54-0400

The amount of heat is


"Q=cm(t_2-t_1)=\\=4.18\cdot84.0\cdot(62.0-12.4)=17716\text{ J}."